# Finte-size induced random switching of chimeras in a deterministic two-population Kuramoto-Sakaguchi model

### v1

@@ V Stochastic model reduction
The numerical results in Section III and (19) suggest to model $Z$ by a Gaussian process $\zeta_{t}$ as in (19). We approximate $\zeta_{t}$ by an Ornstein-Uhlenbeck (OU) process $\hat{\zeta}(t)$, the parameters of which are chosen to match the mean and the covariance function of the true $Z$. We hence replace the deterministic evolution equation (28) by the stochastic evolution equation
@@
$+\textrm{Im}\left[\left(Kr_{1}+\kappa\left(\langle Z\rangle+\frac{1}{\sqrt{N}}\hat{\zeta}\right)\right)e^{-i(\theta_{i}+\lambda-\psi_{1})}\right],$ (29)
@@
with complex drift and diffusion coefficients $\gamma$ and $\sigma$, respectively, and a complex Wiener process $W(t)$. The mean-value $\langle Z\rangle$ can be either numerically estimated from long-time simulations or approximated by the results from the mean-field theory in Section II.1. For the parameters we consider here with $N=12$, $K=100$, $\kappa=60$ and $\lambda=\pi/2-0.075$ and a standard Gaussian native frequency distribution with mean zero and variance $\sigma_{\omega}^{2}=1$ we obtain $\langle Z\rangle=0.6227+0.1000\,i$ from a long-time simulation of the full two-population KS model (4)-(5) over $10,000$ time units which is nicely reproduced by the reduced deterministic system (25) with $\langle Z\rangle=0.6096+0.0934\,i$. In the following simulations we employ the thermodynamic limit.
@@ V.1 Approximation of fluctuations $\zeta$ by an Ornstein-Uhlenbeck process $\hat{\zeta}$
Any Gaussian process is entirely determined by its mean and its covariance function. It is convenient to consider a $2$-dimensional Wiener process and treat $\zeta$ as a $2$-dimensional column vector with real entries. The complex drift and diffusion coefficients are then represented by $2\times 2$ matrices $\Gamma$ and $\Sigma$, erspectively. To avoid notational confusion we denote this real-valued stochastic process of fluctuations as $\vec{\zeta}$. The covariance function of the full two-population KS model (1)–(2) is defined as
@@
$=\langle\vec{\zeta}(s+t)\vec{\zeta}(s)^{\top}\rangle_{s}$ (31)
and is a real-valued $2\times 2$ matrix. The covariance matrix $R(t)$ can be determined numerically from a long-time simulation. The covariance function of an approximating OU process (30) is given explicitly as
@@
where $\vec{\zeta}_{0}$ are random variables drawn from the density of the $2$-dimensional OU-process (30) and the angular brackets $\langle\cdot\rangle_{\rm{OU}}$ denote the average with respect to the stationary Gaussian density of the OU-process. For simplicity, we impose a symmetry assumption between the real and imaginary parts of the drift and diffusion matrices and set
$\Gamma$ $=\operatorname{Re}(\gamma)\operatorname{I}+\operatorname{Im}(\gamma)\operatorname{J}$ (33)
@@
with identity matrix $\operatorname{I}$ and skew-symmetric matrix $\operatorname{J}$ with $J_{12}=-J_{21}=1$ and $J_{11}=J_{22}=0$, rather than considering more general drift and diffusion matrices as done in Yue and Gottwald (2024). The averages in the covariance function of the OU process (32) can be calculated explicitly, see for example Yue and Gottwald (2024). The aim is to determine the parameters of the OU-process (30), i.e. the coefficients $\gamma$ and $\sigma$, such that the covariance function of the OU process (32) best approximates the actual covariance function (V.1). The covariance function $R(t)$ of the fluctuations of the full two-population KS model exhibits small-scale variations that quickly decay as shown in Figure 11. This suggest to only fit the covariance functions for times larger than $t=t_{\rm{min}}$ after which the fast-scale processes have decayed. To find the approximating OU-process we therefore minimise the objective function
@@
We use the MATLAB function fminsearch to perform the nonlinear least-square optimisation problem for (35). We use $t_{\rm min}=0.5$ and $t_{\rm max}=1$. We initialise $\gamma=\gamma_{0}$ as the eigenvalue of the linearisation of the deterministic approximate mean-field dynamics (IV) around the stationary solution, which we numerically estimated as $\gamma_{0}=0.42-28.98\,i$. We found that the optimisation was less sensitive to the choice of the initial diffusion coefficient $\sigma$; we chose $\sigma$ such that $\sigma^{2}/(2N\operatorname{Re}(\gamma))=0.005$. The numerical minimisation of (35) yields
@@ V.2 A Kramers problem for the mean switching time
The reduced stochastic model (V)–(30) suggests that switching is induced by sufficiently large fluctuations $\zeta$. Figure 7 revealed that such fluctuations from the thermodynamic mean $\langle Z\rangle$ (denoted by the red dot) are those that cross the basin boundary denoted by the black line. An approximation for this basin boundary is a circle with radius $q^{\star}$ which we estimated numerically from the full two-population KS model as $q^{\star}=0.386$ and from our reduced deterministic mean-field model (IV) and its averaged approximation (IV) as $q^{\star}=0.339$. This suggests to compute the mean switching time by determining the average time it takes the OU process (30) to generate fluctuations of size $q^{\star}$ when started at equilibrium with $q=0$. This is the classic setting of a Kramers problem Kramers (1940) for a stochastic differential equation for the azimuthally averaged fluctuation
@@
$dq$ $=(\frac{\sigma^{2}}{2qN}-\text{Re}(\gamma)q)\,dt+\frac{\sigma}{\sqrt{N}}dB_{t}$
@@
with the generator associated with the stochastic process (IV)
@@
and boundary conditions $\partial\tau_{e}/\partial q(q=0)=0$ and $\tau_{e}(q^{\star})=0$. The solution of the Kramers problem (41) is explicitly written as
@@
which can be evaluated numerically. Note that the exit time (and hence the mean switching time) is determined by the stochastic dynamics of the fluctuations (IV). Information about the synchronised population only enters via the value of $q^{\star}$.
In the next Section we show that the reduced stochastic model (V)–(30) is able to quantitatively capture the statistical behaviour of the deterministic two-population KS model (4)–(5), including the mean switching time.
@@ VI Comparison of the reduced stochastic model with the full two-population Kuramoto-Sakaguchi model
We now assess the ability of the reduced stochastic model of the synchronised population (V)–(30) to describe the dynamics of the full two-population KS model (1)–(2). We begin by first assessing how the model is able to describe the invariant density of the order parameter $r_{1}$ and of the relative phases $\theta_{i}-\psi_{1}$. The reduced stochastic model (V)–(30) is integrated with a time step of $\Delta t=0.001$ using a 4th order Runge-Kutta scheme with pregenerated noise $\hat{\zeta}$. To generate the OU process $\hat{\zeta}$ we employ a numerical scheme that exactly reproduces the transition probabilities independent of the employed time step (see Appendix B for details).
We first show in Figure 12 a comparison of the empirical histogram of the order parameter $r_{1}$ for the synchronised population for $N=8,12,16$ and $N=32$. Remarkably, even $N=8$ oscillators are sufficient to allow for the reduced stochastic model to reliably approximate the density of the order parameter.
@@
The empirical histograms shown in Figure 12 have pronounced peaks. These peaks are caused by one or more oscillators of the synchronised population to desynchronise, which causes isolated drops of the order parameter as illustrated in Figure 13, which shows the order parameter $r_{1}$ as a function of time, during periods over which no switching occurred. The amplitude of the peaks can be explicitely estimated via a simple argument: Under the assumption that a single oscillator, which we label here without loss of generality as $\theta_{N}$, desynchronises with $\theta_{N}=\pi$ from the synchronised population, which we approximate by $\theta_{1}=\theta_{2}=\dots\theta_{N-1}=0$, the order parameter can be approximated as $r_{1}\approx 1-\frac{2}{N}$, which fits the observed peaks in the histograms depicted in Figure 12 very well, in particular for $N>12$. The frequency of the dips and their size are well recovered by the reduced stochastic model for $N\geq 12$.
@@
The reduced stochastic model (V)–(30) is further able to reproduce the statistics of the fluctuations of the synchronised phases around their mean, $\theta_{i}-\psi_{1}$. Figure 14 shows the empirical histograms of $\theta_{i}-\psi_{1}$ for $N=8,12,16$ and $N=32$, over all oscillators of the synchronised population. It is seen again that even for the small system size with $N=8$ the reduced stochastic model is able to describe the statistical properties of the full two-population KS model (1)–(2). The accuracy of the reduced stochastic model increases, as expected, with increasing system size $N$. The histograms of the full two-population KS model exhibit isolated sharp peaks for $N<16$ at the mean of the phases ofvindividual oscillators. These undulations get smeared out for a sufficiently large number of oscillators.
@@
To see in how far the reduced stochastic model captures the statistics of individual oscillators we show in Figures 15 and 16 the mean and variance for each synchronised oscillator, and in Figure 17 the average effective frequency ${\bar{\nu}}_{i}\coloneq\langle\theta_{i}(t)\rangle$ for each synchronised oscillator. Remarkably, the mean of the fluctuations is well approximated by the reduced stochastic model even for $N=8$. These figures show that the synchronised population involves rogue oscillators at the edge of the population, which oscillate with markedly different mean rotation frequencies and mean phases. The reduced stochastic model is able to capture the mean phase as well as the variance of all oscillators, even of the rogue oscillators very well.
@@
Figure 18 shows the mean switching time $\bar{\tau}$ as a function of the system size $N$. We compare the mean switching time of the full two-population KS model (1)–(2) with the expected switching times $\bar{\tau}=2\tau_{e}$ of the reduced stochastic model (V)–(30) which we estimate from the solution of Kramers problem (43) for both $q^{\star}=0.386$ as estimated from the full two-population KS model (1)–(2) (cf. Figure 7) and for $q^{\star}=0.339$ as estimated from the reduced deterministic mean-field model (cf. Figures 9 and 10). For each value of $N$ the mean switching time is estimated over $1,000$ switching events in the full model. It is remarkable how the simple approximation captures the mean switching time reasonably well. Setting a uniform critical fluctuation size $q^{\star}$ by azimuthally averaging is a significant simplification (which clearly fails at the top right of the blue tear drop depicted in Figure 7). We have checked that for parameter values which exhibit a less circular basin boundary the approximation of the mean switching times becomes worse.
@@
Figure 18: Comparison of the mean switching time $\bar{\tau}$ for different system sizes NN, obtained from a long-time simulations of the full two-population KS model (1)–(2) (blue circles) and from the analytical expressions (40) and (43) based on solving Kramers problem using our reduced stochastic model. We show the result of the Kramers problem for $q^{\star}=0.386$ (orange line) which was estimated from the full two-population KS model (1)–(2) (cf. Figure 7), and for $q^{\star}=0.339$ (yellow line) which was estimated from the reduced deterministic mean-field model (IV) (cf. Figures 9 and 10). Equation parameters are as in Figure 1.
@@ VII Summary and Outlook
We have shown that a deterministic two-population Kuramoto-Sakaguchi model can exhibit switching of chimera states whereby the two populations switch between a synchronised and desynchronised state in a seemingly random way. The switching events were shown to be uncorrelated events and to follow a Poisson distribution. The average time between switches was shown to increase exponentially with the system size $N$ and hence switching constitutes a finite-size effect. As such, the effective stochastic behaviour cannot be described by the standard techniques valid in the thermodynamic limit. We evoked the central limit theorem to describe the fluctuations around the mean behaviour and employed the stochastic modelling framework from Yue and Gottwald (2024). The effective reduced stochastic model captures the statistics of both, the synchronised cluster and the switching of this cluster into a desynchronised state. Using mean-field theory we obtained an approximation of the critical size of finite-size fluctuations required to induce chimera switching. Our model produces relatively accurate predictions for the rate at which switching events become less frequent with increasing system size.
For illustration purposes, we have presented results for a single set of parameters. We present in Appendix C further results for a different set of parameters to show the robustness of our approach.
@@
Our approach here is entirely heuristic and relies on numerically fitting an Ornstein-Uhlenbeck process to the observed statistics of the fluctuations. It would be interesting to see if the approximation of the OU process can at least be formally justified. It was argued that the approximation of $Z$ is related to a trigonometric approximation of Gaussian processes, where instead of trigonometric functions, the near-periodic solutions of the desynchronised oscillators are used Yue and Gottwald (2024). This suggests a theoretical avenue to put the heuristic method employed here on a more rigorous footing.
@@
The $N$-dimensional reduced stochastic model is still relatively high-dimensional and it would be interesting to see if it can be further reduced. Since the system-size $N$ is not sufficiently large to allow for mean-field theory to capture the finite-size induced switching we need to resort to methods which are valid for finite-size systems such as the method of collective coordinates Gottwald (2015); Hancock and Gottwald (2018); Smith and Gottwald (2019, 2020); Yue, Smith, and Gottwald (2020) which has already been shown to work for stochastically driven Kuramoto models Gottwald (2017). This would allow for a straightforward calculation of the Kramers problem to calculate the transition probabilities describing switching chimeras.

### v3

@@ V Stochastic model reduction
The numerical results in Section III suggest modelling $Z$ by a Gaussian process $\zeta_{t}$ as in (22). We approximate $\zeta_{t}$ by an Ornstein-Uhlenbeck (OU) process $\hat{\zeta}(t)$, the parameters of which are chosen to match the mean and the covariance function of the true $Z$. We hence replace the deterministic evolution equation (30) by the stochastic evolution equation
@@
$+\textrm{Im}\left[\left(Kr_{1}+\kappa\left(\langle Z\rangle+\frac{1}{\sqrt{N}}\hat{\zeta}\right)\right)e^{-i(\theta_{i}+\lambda-\psi_{1})}\right]$ (31)
@@
with complex drift and diffusion coefficients $\gamma$ and $\sigma$, respectively, and a complex-valued Wiener process $W(t)$. The mean $\langle Z\rangle$ can be either estimated numerically from a long-time simulation or by the mean-field theory in Section II.1. For the parameters we consider here with $N=12$, $K=100$, $\kappa=60$ and $\lambda=\pi/2-0.075$ and a standard Gaussian native frequency distribution with mean zero and variance $\sigma_{\omega}^{2}=1$ we obtain $\langle Z\rangle=0.6227+0.1000\,i$ from a long-time simulation of the full two-population KS model (5)–(6) over $10,000$ time units which is nicely reproduced by the reduced deterministic system (27) with $\langle Z\rangle=0.6096+0.0934\,i$. In the following simulations, we use the thermodynamic limit obtained from (27).
@@ V.1 Approximation of fluctuations $\zeta$ by an Ornstein-Uhlenbeck process $\hat{\zeta}$
Any Gaussian process is entirely determined by its mean and its covariance function. It is convenient to represent the complex-valued fluctuations $\zeta$ by a $2$-dimensional real-valued Wiener process and treat $\zeta$ as a $2$-dimensional column vector with real entries. The complex drift and diffusion coefficients are then represented by $2\times 2$ real matrices $\Gamma$ and $\Sigma$, respectively. To avoid notational confusion, we denote this real-valued stochastic process of fluctuations as $\vec{\zeta}$. The covariance function of the full two-population KS model (1)–(2) is defined as
@@
$=\langle\vec{\zeta}(s+t)\vec{\zeta}^{\top}(s)\rangle_{s}.$ (33)
The covariance matrix $R(t)$ can be determined numerically from a long-time simulation. The covariance function of an approximating OU process (32) is given explicitly as
@@
where $\vec{\zeta}_{0}$ are random variables drawn from the density of the $2$-dimensional OU process (32) and the angular brackets $\langle\cdot\rangle_{\rm{OU}}$ denote the average with respect to the stationary Gaussian density of the OU process. For simplicity, we impose a symmetry assumption between the real and imaginary parts of the drift and diffusion matrices and set
$\Gamma$ $=\operatorname{Re}(\gamma)\operatorname{I}+\operatorname{Im}(\gamma)\operatorname{J},$ (35)
@@
with identity matrix $\operatorname{I}$ and skew-symmetric matrix $\operatorname{J}$ with $J_{12}=-J_{21}=1$ and $J_{11}=J_{22}=0$, rather than considering more general drift and diffusion matrices as done in Yue and Gottwald (2024). The averages in the covariance function of the OU process (34) can be calculated explicitly, see for example Yue and Gottwald (2024). The aim is to determine the parameters of the OU process (32), i.e. the coefficients $\gamma$ and $\sigma$, such that the covariance function of the OU process (34) best approximates the actual covariance function (V.1). The covariance function $R(t)$ of the fluctuations of the full two-population KS model exhibits small-scale variations that quickly decay, as shown in Figure 10. This suggests only fitting the covariance functions for times larger than $t=t_{\rm{min}}$ after which the fast-scale processes have decayed. To find the approximating OU process, we therefore minimise the objective function
@@
We use the MATLAB function fminsearch to perform the nonlinear least-square optimisation problem for (37). We use $t_{\rm min}=0.5$ and $t_{\rm max}=1$ and initialise $\gamma=\gamma_{0}$ as the eigenvalue of the linearisation of the deterministic approximate mean-field dynamics (IV) around the stationary solution, which we numerically estimated as $\gamma_{0}=0.42-28.98\,i$. We found that the optimisation was less sensitive to the choice of the initial diffusion coefficient $\sigma$; we chose $\sigma$ such that $\sigma^{2}/(2N\operatorname{Re}(\gamma))=0.005$. The numerical minimisation of (37) yields
@@ V.2 A Kramers problem for the mean switching time
The reduced stochastic model (V)–(32) suggests that switching is induced by sufficiently large fluctuations $\zeta$. Figure 6 suggests that such fluctuations from the thermodynamic mean $\langle Z\rangle$ (denoted by the red dot) are those that cross the basin boundary denoted by the black line. An approximation for this basin boundary is a circle with radius $q^{\star}$, which we estimated numerically from the full two-population KS model as $q^{\star}=0.386$ and from our reduced deterministic mean-field model (IV) and its averaged approximation (29) as $q^{\star}=0.339$. This suggests computing the mean switching time by determining the average time it takes the OU process (32) to generate fluctuations of magnitude $q^{\star}$ when started at equilibrium with $q=0$. This is the classic setting of a Kramers problem Kramers (1940) for a stochastic differential equation for the azimuthally averaged fluctuation
@@
$dq$ $=\left(\frac{\sigma^{2}}{2qN}-\text{Re}(\gamma)q\right)\,dt+\frac{\sigma}{\sqrt{N}}dB_{t}$
@@
with the generator associated with the stochastic process (29)
@@
and boundary conditions $\partial\tau_{e}/\partial q(q=0)=0$ and $\tau_{e}(q=q^{\star})=0$. The solution of the Kramers problem (43) is explicitly written as
@@
which can be evaluated numerically. Note that the exit time (and hence the mean switching time) is determined by the stochastic dynamics of the fluctuations (29). Information about the synchronised population only enters via the value of $q^{\star}$.
In the next section, we show that the reduced stochastic model (V)–(32) is able to quantitatively capture the statistical behaviour of the deterministic two-population KS model (5)–(6), including the mean switching time.
@@ VI Comparison of the reduced stochastic model with the full two-population Kuramoto-Sakaguchi model
We now assess the ability of the reduced stochastic model of the synchronised population (V)–(32) to describe the dynamics of the full two-population KS model (1)–(2). We begin by first assessing how the model is able to describe the invariant density of the order parameter $r_{1}$ and of the relative phases $\theta_{i}-\psi_{1}$ of individual synchronised oscillators. The reduced stochastic model (V)–(32) is integrated with a time step of $\Delta t=0.001$ using a 4th order Runge-Kutta scheme with pregenerated noise $\hat{\zeta}$. We remark that this scheme is not fourth-order in time due to the presence of the non-smooth noise. To generate the OU process $\hat{\zeta}$ we employ a numerical scheme that exactly reproduces the transition probabilities independent of the employed time step (see Appendix B for details).
@@
We first show in Figure 11 a comparison of the empirical histogram of the order parameter $r_{1}$ for the synchronised population for $N=8,12,16$ and $N=32$. Remarkably, even $N=8$ oscillators are sufficient to allow for the reduced stochastic model to accurately approximate the density of the order parameter. The empirical histograms shown in Figure 11 have pronounced peaks. These peaks are caused by one or more oscillators of the synchronised population to phase-slip, which causes isolated drops of the order parameter as illustrated in Figure 12, which shows the order parameter $r_{1}$ as a function of time, during periods over which no switching occurred. The frequency of the dips and their size are well recovered by the reduced stochastic model for $N\geq 12$. The amplitude of the peaks can be explicitly estimated via a simple argument: Under the assumption that a single oscillator, which we label here without loss of generality as $\theta_{N}$, desynchronises with $\theta_{N}=\pi$ from the synchronised population, which we approximate by $\theta_{1}=\theta_{2}=\dots\theta_{N-1}=0$, the order parameter can be approximated as $r_{1}\approx 1-\frac{2}{N}$, which fits the observed peaks in the histograms in Figure 11 and in the time series in Figure 12 very well, in particular for $N>12$.
@@
The reduced stochastic model (V)–(32) further reproduces the statistics of fluctuations of the synchronised phases around their mean, $\theta_{i}-\psi_{1}$. Figure 13 shows the empirical histograms of $\theta_{i}-\psi_{1}$ for $N=8,12,16$ and $N=32$, over all oscillators of the synchronised population. It is seen again that even for a small system size with $N=8$ the reduced stochastic model is able to describe the statistical properties of the full two-population KS model (1)–(2). The accuracy of the reduced stochastic model increases, as expected, with increasing system size $N$. The histograms of the full two-population KS model exhibit isolated sharp peaks for $N<16$ at the mean of the phases of individual oscillators. These undulations get smeared out for a sufficiently large number of oscillators.
@@
To see in how far the reduced stochastic model captures the statistics of individual oscillators we show in Figures 14 and 15 the mean and variance for each synchronised oscillator, and in Figure 16 the average effective frequency ${\bar{\nu}}_{i}\coloneq\langle\theta_{i}(t)\rangle$ for each synchronised oscillator. Remarkably, the mean of the fluctuations is well approximated by the reduced stochastic model even for $N=8$. These figures show that the synchronised population involves rogue oscillators at the edge of the population, which oscillate with markedly different mean rotation frequencies and mean phases. The reduced stochastic model is able to capture the mean phase as well as the variance of all oscillators very well, even for the rogue oscillators.
@@
The empirical histograms of the order parameter and the statistics of the fluctuations of individual synchronised oscillators are relatively robust to the numerically estimated covariance function. In contrast, rare switching events are induced by large excursions of the fluctuations and hence depend crucially on correctly estimating the tail behaviour of their distribution. Figure 17 shows the mean switching time $\bar{\tau}$ as a function of the system size $N$. We compare the mean switching time of the full two-population KS model (1)–(2) with the expected switching times $\bar{\tau}=2\tau_{e}$ of the reduced stochastic model (V)–(32) which we estimate from the solution of the Kramers problem (45) for both $q^{\star}=0.386$ as estimated from the full two-population KS model (1)–(2) (cf. Figure 6) and for $q^{\star}=0.339$ as estimated from the reduced deterministic mean-field model (cf. Figures 8 and 9). For each value of $N$, the mean switching time is estimated over $1,000$ switching events in the full model. It is remarkable how our simple approximation captures the mean switching time reasonably well. Setting a uniform critical fluctuation size $q^{\star}$ by azimuthally averaging is a significant simplification (which clearly fails at the top right of the blue teardrop depicted in Figure 6). We have checked that for parameter values that exhibit a less circular basin boundary, the approximation of the mean switching times becomes worse.
@@
Figure 17: Comparison of the mean switching time $\bar{\tau}$ for different system sizes NN, obtained from long-time simulations of the full two-population KS model (1)–(2) (blue circles) and from the analytical expressions (42) and (45) based on solving Kramers problem using our reduced stochastic model. We show results of the Kramers problem for $q^{\star}=0.386$ (red line), which was estimated from the full two-population KS model (1)–(2) (cf. Figure 6), and for $q^{\star}=0.339$ (yellow line), which was estimated from the reduced deterministic mean-field model (IV) (cf. Figures 8 and 9). Equation parameters are as in Figure 1.
@@ VII Summary and Outlook
We have shown that a deterministic two-population Kuramoto-Sakaguchi model exhibits switching of chimera states whereby the two populations switch between a synchronised and desynchronised state in a seemingly random way. The switching events were shown to be uncorrelated events and to follow a Poisson distribution. The average time between switches was shown to increase exponentially with the system size $N$ and hence switching constitutes a finite-size effect. As such, the effective stochastic behaviour cannot be described by the standard techniques valid in the thermodynamic limit. We evoked the central limit theorem to describe the fluctuations around the mean behaviour and employed the stochastic modelling framework from Yue and Gottwald (2024). The effective reduced stochastic model captures the statistics of both the synchronised cluster and the switching of this cluster into a desynchronised state. Using mean-field theory, we obtained an approximation of the critical size of finite-size fluctuations required to induce chimera switching. Our model produces relatively accurate predictions for the rate at which switching events become less frequent with increasing system size.
The chimera switching in the two-population KS model is a non-transient persistent process with a stationary probability density. in contrast, in a single population KS model with identical oscillators, chimera states were shown to be neutrally stable in the thermodynamic limit and were in fact only chaotic transients in finite-size systems, collapsing to a stable synchronised state WolfrumOmelchenko11 ; WolfrumEtAl11 . In our two-population system, the state where both populations are synchronised is not stable. This, we conjecture, leads to the emergence of persistent chimera switching. To fully analyse this mechanism is planned for further research.
For illustrative purposes, we have presented results for a single set of parameters and with equiprobably drawn native frequencies. Our results are robust with respect to changes in the parameters (see Appendix C). For random realisations of native frequency draws, the probability of observing switching in a finite time window sensitively depends on the details of the realisation. In particular, we find that switching is suppressed if at least one of the populations has a smaller than expected variance of their native frequencies; this can be understood by realising that tighter native frequency distributions promote stability of the synchronised state. Conversely, switching is promoted by an overall larger spread of the native frequencies. This, we argue, in turn increases the tail in the distribution of the fluctuations $\zeta$, responsible for inducing switching (see Appendix D, for details).
@@
Our approach here is entirely heuristic and relies on numerically fitting an Ornstein-Uhlenbeck process to the observed statistics of the fluctuations. It would be interesting to see if the approximation of the OU process can at least be formally justified. It was argued that the approximation of $Z$ is related to a trigonometric approximation of Gaussian processes, where instead of trigonometric functions, the near-periodic solutions of the desynchronised oscillators are used Yue and Gottwald (2024). This suggests a theoretical avenue to put the heuristic method employed here on a more rigorous footing, following the recent work of OmelchenkoGottwald25 .
@@
The $N+2$-dimensional reduced stochastic model is still relatively high-dimensional, and it would be interesting to see if it can be further reduced. Since the system-size $N$ is not sufficiently large to allow for mean-field theory to capture the finite-size induced switching we need to resort to methods which are valid for finite-size systems such as the method of collective coordinates Gottwald (2015); Hancock and Gottwald (2018); Smith and Gottwald (2019, 2020); Yue, Smith, and Gottwald (2020) which has already been shown to work for stochastically driven Kuramoto models Gottwald (2017). This would allow for a straightforward calculation of the Kramers problem to calculate the transition probabilities describing switching chimeras.